OpenStudy (anonymous):

Drew put $4500 in an account with an annual simple interest rate of 5%. When he closed the account, the balance was $5850. For how long did Drew have the account?

jimthompson5910 (jim_thompson5910):

A = P*(1+r*t) 5850 = 4500*(1+0.05*t) Now solve for t. Tell me what you get.

OpenStudy (anonymous):

i got 26..

jimthompson5910 (jim_thompson5910):

that's way too high 5850 = 4500*(1+0.05*t) 5850/4500 = 1+0.05*t 1.3 = 1+0.05*t 1 + 0.05t = 1.3 0.05t = 1.3 - 1 .... .... ... t = ???

OpenStudy (anonymous):

i think im gonna need another formula i dont get this one :-/

jimthompson5910 (jim_thompson5910):

what is 1.3 - 1 equal to

OpenStudy (anonymous):

0.3

jimthompson5910 (jim_thompson5910):

so 0.05t = 0.3

OpenStudy (anonymous):

what is the next step

jimthompson5910 (jim_thompson5910):

to undo that multiplication of 0.05

jimthompson5910 (jim_thompson5910):

so what undoes multiplication?

OpenStudy (anonymous):

so divide 0.05/ 0.3 ??

jimthompson5910 (jim_thompson5910):

close, the other way around 0.3/0.05

OpenStudy (anonymous):

ohh ok

jimthompson5910 (jim_thompson5910):

0.05t = 0.3 0.05t/0.05 = 0.3/0.05 t = ???

OpenStudy (anonymous):

6..

OpenStudy (anonymous):

so is the answer gonna be 6 years?

jimthompson5910 (jim_thompson5910):

yes, so t = 6 which means it takes 6 years

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

ok ok thank u

jimthompson5910 (jim_thompson5910):

you're welcome